### v1

# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High Dimensional Parameters

Abstract
This paper extends the Nested Error Regression Model with High-Dimensional Parameters (NERHDP) to address challenges in small area poverty estimation. Building on the NERHDP framework, we develop a robust and flexible approach to derive empirical best predictors (EBPs) of small area poverty indicators, while accommodating heterogeneity in regression coefficients and sampling variances across areas. To overcome computational limitations in the existing algorithm, we introduce an efficient method that significantly reduces computation time, enhancing scalability for large datasets. Additionally, we propose a novel method for generating area-specific poverty estimates for out-of-sample areas, improving the reliability of synthetic estimates. To quantify uncertainty, we introduce a parametric bootstrap method tailored to the extended model. Through design-based simulation studies, we demonstrate that the proposed method has better performance in terms of relative bias and relative root mean squared prediction error compared to existing approaches. Furthermore, the proposed method is applied to household survey data from the 2002 Albania Living Standards Measurement Survey to estimate poverty indicators for 374 municipalities using auxiliary information from the 2001 census.
Poverty mapping,
small area estimation,
data linkage,
robust inference,
keywords:
[…]
, and
1 Introduction
The United Nations emphasizes the priority of eradicating all forms of poverty by making it the first of its 17 Sustainable Development Goals. Measuring poverty levels across regions is essential for designing effective poverty reduction strategies and ensuring equitable allocation of resources. Accurate information about people’s living conditions at a regional level is a basic instrument for targeting policies and programs to reduce poverty. However, progress toward this goal is often slowed by a consistent lack of data on important economic indicators, especially for small population groups and specific geographic areas in developing countries. One major reason for the shortage of of disaggregated poverty data is that large scale sample surveys are typically designed to produce reliable estimates of various characteristics of interest for large geographic areas or large population subgroups, rather than for smaller, more granular domains (Chatterjee, Lahiri and Li, 2008).
Standard survey estimates for national and major subgroups are referred to as ‘design-based’ or ‘direct’ estimates, as they rely solely on the survey data and the selection probabilities specific to the subgroup of interest. The direct (or design-based) estimates become problematic when the sample sizes in the subgroups of interest are small (or even zero). In this situation, model-based methods are increasingly being used to produce what are termed as ‘small area’ or ‘indirect’ estimates, which ‘borrow strength’ across areas through linking models based on auxiliary information coming from censuses or administrative registers; for a detailed overview of these techniques, refer to Jiang and Nguyen (2007) and Rao and Molina (2015).
Many measures of poverty and inequality are nonlinear functions of the small area distribution of the economic variable underpinning poverty, usually income or a surrogate such as expenditure. In this paper, we focus on estimating a class of poverty measures known as the Foster–Greer–Thorbecke (FGT) poverty measures, introduced by Foster, Greer and Thorbecke (1984). Consider a finite population of size $N$ partitioned into $m$ small areas of size $N_{1},\cdots,N_{m}$. Let $E_{ij}$ be a suitable quantitative measure of welfare for individual $j$ in small area $i$, such as income or expenditure, and let $z$ be a fixed poverty line. This poverty line $z$ serves as the threshold below which an individual is considered to be living in poverty. The FGT family of poverty measures for each small area $i$ is then defined as the area-specific mean
$F_{\alpha i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}F_{\alpha ij},\;i=1,\cdots,m,$ (1)
of the values $F_{\alpha ij}$ defined as
$F_{\alpha ij}=\bigg{(}\frac{z-E_{ij}}{z}\bigg{)}^{\alpha}I(E_{ij}<z),\;j=1,\cdots,N_{i};\;\alpha=0,1,2,$ (2)
where $I(E_{ij}<z)=1$ if $E_{ij}<z$, otherwise, it equals to 0. When $\alpha=0$, the FGT measure represents the proportion of individuals living below the poverty line in small area $i$, commonly referred to as the poverty incidence or headcount ratio. For $\alpha=1$, the measure, known as the poverty gap, captures the average relative distance of individuals to the poverty line within the area. When $\alpha=2$, the measure, termed poverty severity, highlights the inequality among the poor, with higher values of $F_{2i}$ indicating areas experiencing more severe levels of poverty.
The small area estimation (SAE) literature for poverty measures was revolutionized by the pioneering work of Elbers, Lanjouw and Lanjouw (2003), which introduced a unit-level approach (ELL) for poverty mapping. This method has become widely adopted for producing fine-grained estimates of FGT measures. The ELL approach assumes a two-level model with a common regression relationship between the response and auxiliary variables while allowing for heteroskedastic unit-level errors. The unit-level variance components are estimated using a proposed alpha model. Simulated census data are then generated to estimate poverty measures. It is important to note that the ELL approach is a synthetic method that does not rely on a conditional distribution. For further details, we refer readers to their original paper.
Another popular approach for poverty estimation introduced by Molina and Rao (2010) is the empirical best prediction (EBP) method through Monte Carlo simulations. This method assumes that the transformed population welfare variables is characterized by a homoskedastic nested error regression (NER) model, as proposed by Battese, Harter and Fuller (1988). Specifically, it assumes that the regression coefficients and variance components are the same across all small areas, with normally distributed errors. Model parameters are estimated using the maximum likelihood (ML) method or the restricted maximum likelihood (REML) method, and EBPs of poverty indicators are then approximated through random draws from the conditional distribution of the out-of-sample units, given the sample data. Molina, Nandram and Rao (2014) introduced a Bayesian version of this approach, which is particularly useful for quantifying the uncertainty of the estimates.
The effectiveness of such EBP method tends to decline when the number of small areas to be combined becomes large. This limitation arises because the assumption of identical regression coefficients and/or variance components across all small areas, inherent in these models, may not hold universally, leading to potential issues with model misspecification. To address this, random area-specific regression coefficient models (Prasad and Rao, 1990; Hobza and Morales, 2013; Rao and Molina, 2015) and random area-specific sampling variance models (Otto and Bell, 1995; Arora, Lahiri and Mukherjee, 1997; Kubokawa et al., 2016; Sugasawa and Kubokawa, 2017) have been suggested in the literature. Such modeling, though useful in some applications, needs more nontrivial assumptions on the joint distribution for a large number of random effects. On the other hand, fixed effects assumptions on the area specific regression coefficients and sampling variances generally lead to unstable estimates of these fixed effects due to small area specific sample sizes (Jiang and Nguyen, 2012).
Lahiri and Salvati (2023) introduced a novel approach in the SAE literature, addressing several challenges. Specifically, this approach assumes fixed effects for both area-specific regression coefficients and sampling variances. But they estimate these coefficients using area-specific (with respect to a tuning parameter) estimating equations applied to data from all areas, and then use appropriately constructed residuals for estimation of variance components. When the tuning parameters are known, they haven shown that the model parameters can be consistently estimated. This approach, termed as the Nested Error Regression Model with High-Dimensional Parameters (NERHDP), not only offers a more flexible SAE modeling framework but also provides a reliable solution to the challenges of high-dimensional parameter estimation in small area analysis. However, their study focused solely on empirical best predictors (EBPs) for small area means, which are linear functions of the data.
In this paper, we extend the NERHDP model to obtain EBPs of poverty indicators, as particular complex and nonlinear FGT measures. Unlike the approach in Molina and Rao (2010), our proposed EBPs account for heterogeneity across small areas in both regression coefficients and sampling variances. Such heterogeneity is common in real-world applications due to variations in socio-economic conditions, data quality, and sampling designs among areas. By extending NERHDP to accommodate FGT poverty measures, we enable the generation of more precise and reliable estimates that are sensitive to area-specific characteristics. This innovation not only enhances the flexibility and applicability of the modeling but also provides a robust methodological framework for addressing the complexities inherent in small area poverty estimation.
Additionally, we address the computational challenges associated with the algorithm proposed by Lahiri and Salvati (2023), which involves iterative convergence and can be computationally intensive. We introduce a more efficient algorithm that significantly reduces computation time, producing estimates within seconds. This improvement ensures the proposed model is not only methodologically robust but also practical for large-scale applications.
Another key feature of our extension is its ability to handle out-of-sample areas, which is particularly important when analyzing a large number of areas. In Molina and Rao (2010), the EBPs become purely synthetic estimates for non-sampled areas due to the absence of observations. In this study, we introduce a new method that enables the estimation of area-specific model parameters even for non-sampled areas. This approach allows for the estimation of poverty measures that are more tailored to specific areas and so less synthetic. While these estimates remain synthetic, they better capture area-specific heterogeneity compared to those in Molina and Rao (2010). Finally, we propose a parametric bootstrap method for quantifying the uncertainty associated with the estimated poverty measures.
We conduct design-based simulation studies using the 2021 Italian dataset from the European Survey on Income and Living Conditions as the fixed population, which exhibits potential heterogeneity in both regression coefficients and sampling variances, to evaluate the performance of our proposed method. We compare it with existing approaches, including the EBP method of Molina and Rao (2010) (MR), the synthetic method of Elbers, Lanjouw and Lanjouw (2003) (ELL), and the direct estimator. The simulation results demonstrate that our method has better performance than the existing approaches in estimating headcount ratios and poverty gaps for both in-sample and out-of-sample areas, based on relative bias (RB) and relative root mean squared prediction error (RRMSPE).
We demonstrate the benefits of implementing the proposed approach on data from the 2002 Living Standards Measurement Survey (LSMS) in Albania to estimate headcount ratios and poverty gaps for the 374 municipalities, using auxiliary information from the 2001 Census. Of these municipalities, $161$ are out-of-sample, meaning they were not included in the original probability sample used for estimation. Consequently, their characteristics were not directly observed in the survey, and any estimates for these areas rely entirely on model-based predictions.
The structure of the paper is as follows. Section 2 provides a brief review of the empirical best prediction (EBP) methodology for FGT measures. Section 3 introduces the 2002 Living Standards Measurement Survey (LSMS) in Albania, describes the estimation problem, and outlines the available auxiliary information. Section 4 introduces the best prediction approach under a nested error linear regression model with high-dimensional parameters, along with the associated new parameter estimation algorithms. In Section 5, we describe the parametric bootstrap method used to estimate measures of uncertainty. Section 6 presents the results of design-based simulation studies, comparing the performance of the proposed method with existing approaches in terms of bias and variability. In Section 7, we apply the new methodology to LSMS data to estimate headcount ratios (HCRs) and poverty gaps (PGs) at the municipality level. Finally, Section 8 concludes the paper by summarizing the key findings and suggesting directions for future research.
2 A review of Empirical Best Prediction for FGT measures
Like in Molina and Rao (2010), we assume that a random sample of size $n<N$ is drawn from the finite population using a specified sampling design. Let $\Omega$, $s\subset\Omega$, and $r=\Omega-s$ denote the set of indices for all population units, the set of sampled units, and the set of indices for the unsampled units with size $N-n$, respectively. For area $i$, we have the following associated notations: $\Omega_{i}$, $s_{i}$, and $n_{i}$. Note that if area $i$ is not sampled, then $n_{i}=0$.
Following the notations from Molina and Rao (2010), we consider a random vector, $\mathbf{y}=(Y_{1},\cdots,Y_{N})^{\prime}$, denoting the values of a random variable associated with the $N$ units of a finite population. Let $\mathbf{y}_{s}$ be the sub-vector of $\mathbf{y}$ corresponding to sample elements $s$ and $\mathbf{y}_{r}$ be the sub-vector of out-of sample elements $r$. After reordering the units of the population, we can write $\mathbf{y}=(\mathbf{y}_{s}^{\prime},\mathbf{y}_{r}^{\prime})^{\prime}$. The goal is to predict the value of a real-valued function $\delta=h(\mathbf{y})$ of the random vector $\mathbf{y}$ using the sample data $\mathbf{y}_{s}$. The best predictor (BP) of $\delta$, which minimizing the mean squared prediction error (MSPE), is given by the conditional expectation
$\hat{\delta}^{B}=\text{E}_{\mathbf{y}_{r}}(\delta|\mathbf{y}_{s}),$ (3)
where the expectation is taken with respect to the conditional distribution of $\mathbf{y_{r}}$ given the observed sample data $\mathbf{y}_{s}$.
Suppose there is a one-to-one transformation $Y_{ij}=T(E_{ij})$ of the welfare variables, $E_{ij}$. The vector $\mathbf{y}$ contains the transformed variables $Y_{ij}$ for all population units, following a normal distribution with mean $\bm{\mu}$ and variance $\mathbf{V}$. Using this transformation, the random variable $F_{\alpha ij}$ defined in (2) can be rewritten as:
$F_{\alpha ij}=\bigg{(}\frac{z-T^{-1}(Y_{ij})}{z}\bigg{)}^{\alpha}I(T^{-1}(Y_{ij})<z)=:h_{\alpha}(Y_{ij}),\;j=1,\cdots,N_{i}.$
From (3), the BP of $\delta=F_{\alpha i}$ is given by:
$\hat{F}^{B}_{\alpha i}=\text{E}_{\mathbf{y}_{r}}(F_{\alpha i}|\mathbf{y}_{s})=\frac{1}{N_{i}}\left\{\sum_{j\in s_{i}}F_{\alpha ij}+\sum_{j\in r_{i}}\hat{F}^{B}_{\alpha ij}\right\},$ (4)
where $s_{i}$ and $r_{i}$ denote the sets of sample and out-of-sample units belonging to area $i$, respectively, and $\hat{F}^{B}_{\alpha ij}$ is the BP of $F_{\alpha ij}=h_{\alpha}(Y_{ij})$. The BP of $F_{\alpha ij}$ for $j\in r_{i}$ is given by:
[…]
where $f_{Y_{ij}}(t|\mathbf{y}_{s})$ denotes the conditional density of $Y_{ij}$ given the observed sample data $\mathbf{y}_{s}$.
Due to the complexity of $h_{\alpha}(t)$, it may not be feasible to explicitly calculate the expectation in (5). Assuming that $\mathbf{y}=(\mathbf{y}_{s}^{\prime},\mathbf{y}_{r}^{\prime})^{\prime}$ follows a multivariate normal distribution with mean vector $\bm{\mu}=(\bm{\mu}_{s}^{\prime},\bm{\mu}_{r}^{\prime})^{\prime}$ and a covariance matrix partitioned as:
$\mathbf{V}=\begin{pmatrix}\mathbf{V}_{s}&\mathbf{V}_{sr}\\ \mathbf{V}_{rs}&\mathbf{V}_{r}\\ \end{pmatrix},$
where $\mathbf{V}_{s}=\text{var}(\mathbf{y}_{s})$, $\mathbf{V}_{r}=\text{var}(\mathbf{y}_{r})$, and $\mathbf{V}_{sr}=\mathbf{V}_{rs}^{\prime}=\text{cov}(\mathbf{y}_{s},\mathbf{y}_{r})$. In the absence of sample selection bias, the conditional distribution of $\mathbf{y}_{r}$ given $\mathbf{y}_{s}$ can be expressed as:
$\mathbf{y}_{r}|\mathbf{y}_{s}\sim N(\bm{\mu}_{r|s},\mathbf{V}_{r|s}),$ (6)
where the conditional mean $\bm{\mu}_{r|s}$ and conditional variance $\mathbf{V}_{r|s}$ are given by:
$\bm{\mu}_{r|s}=\bm{\mu}_{r}+\mathbf{V}_{rs}\mathbf{V}^{-1}_{s}(\mathbf{y}_{s}-\bm{\mu}_{s})\;\text{and}\;\mathbf{V}_{r|s}=\mathbf{V}_{r}-\mathbf{V}_{rs}\mathbf{V}^{-1}_{s}\mathbf{V}_{sr}.$ (7)
Molina and Rao (2010) proposed using an empirical approximation of the conditional expectation in (5) through Monte Carlo simulations. Specifically, a large number $K$ of vectors $\mathbf{y}_{r}$ are generated from the conditional distribution in (6). Let $Y_{ij}^{(k)}$ denote the $k$th simulated value of the out-of-sample observation $Y_{ij}$ for $j\in r_{i}$, where $k=1,\cdots,K$. The Monte Carlo approximation to the BP for $Y_{ij}$ for $j\in r_{i}$ is then expressed as:
$\hat{F}^{B}_{\alpha ij}=\text{E}_{\mathbf{y}_{r}}[h_{\alpha}(Y_{ij})|\mathbf{y}_{s}]\approx\frac{1}{K}\sum_{k=1}^{K}h_{\alpha}(Y_{ij}^{(k)}),\;j\in r_{i}.$ (8)
Since the mean vector $\bm{\mu}$ and covariance matrix $\mathbf{V}$ typically depend on an unknown parameter vector $\bm{\phi}$, the conditional density $f_{Y_{ij}}(t|\mathbf{y}_{s})$ also depends on $\bm{\phi}$. Substituting $\bm{\phi}$ with its estimate $\hat{\bm{\phi}}$, we generate simulated values $Y_{ij}^{(k)}$ from the estimated density $f_{Y_{ij}}(t|\mathbf{y}_{s},\hat{\bm{\phi}})$. The resulting empirical best predictor (EBP) of the poverty measure $F_{\alpha i}$ is:
$\hat{F}^{EB}_{\alpha i}=\frac{1}{N_{i}}\left[\sum_{j\in s_{i}}F_{\alpha ij}+\sum_{j\in r_{i}}\hat{F}^{EB}_{\alpha ij}\right].$
Note that the EBP method from Molina and Rao (2010) requires linking survey and census households to identify sample indices, which may not be feasible in real-world applications. Often, the survey sample is not a subset of the census. To address this limitation, we follow the approach proposed by Rodas, Molina and Nguyen (2021) and apply Monte Carlo simulations to all population units rather than restricting it to non-sampled units. Thus, the EBP can be expressed as:
$\hat{F}^{EB}_{\alpha i}=\frac{1}{N_{i}}\sum_{j}^{N_{i}}\hat{F}^{EB}_{\alpha ij}.$ (9)
For consistency, and without causing confusion, we use (9) as the EBP for the poverty measures throughout this paper. When the sampling fraction $n_{i}/N_{i}$ is negligible, which is the most cases in practice, the EBP (9) is practically equivalent to the original formulation proposed by Molina and Rao (2010) and further refined by Rodas, Molina and Nguyen (2021) and Molina, Corral and Nguyen (2022).
3 The LSMS and Census data
The primary dataset used in this study is the 2002 Albania Living Standards Measurement Survey (LSMS), conducted by INSTAT in the Spring of 2002. This survey provides comprehensive insights into the living conditions of the Albanian population, capturing both income and non-income dimensions of poverty. The LSMS sample consists of $3,591$ households, covering a broad range of socio-economic indicators critical for poverty assessment in Albania.
The sample design follows a two-stage cluster sampling strategy, with $450$ primary sampling units (PSUs) selected from the 2001 pre-census enumeration areas (EAs). Within each PSU, $8$ households were randomly chosen. The sampling frame was stratified into four main regions to ensure representativeness across Tirana, other urban areas, rural areas, and the three major agro-ecological/economic zones (Coastal, Central, and Mountain). The final dataset comprises of:
  • $1,000$ households from the Coastal regions,
  • $1,000$ households from the Mountain regions,
  • $991$ households from the Central region,
  • $600$ households from Tirana.
In the sample, a total of $1,640$ households are rural, and $1,951$ are urban.
The 2002 LSMS data includes observations from 213 of the 374 municipalities in Albania, while the remaining 161 municipalities are entirely out-of-sample, meaning no households were directly surveyed in these areas. Consequently, poverty estimates for these out-of-sample municipalities rely solely on model-based predictions, introducing additional uncertainty and potential extrapolation bias if the characteristics of the out-of-sample areas differ significantly from those of the sampled ones.
Among the observed municipalities, sample sizes vary significantly, ranging from just 6 households in municipalities such as Koder Thu and Hajmel to 600 households in Tirana. The mean sample size is 16.8, with quartiles at 8 (25th percentile), 8 (50th percentile), and 16 (75th percentile). This indicates that several municipalities have very small sample sizes, which hinders the reliability of direct estimates.
For sampled municipalities, we compute direct estimates of FGT poverty measures, using the weighted average estimator:
$\hat{F}_{\alpha i}^{w}=\frac{1}{\hat{N}_{i}}\sum_{j\in s_{i}}w_{ij}F_{\alpha ij},$
where $w_{ij}$ represents the survey weight for individual $j$ in municipality $i$, and $\hat{N}_{i}=\sum_{j\in s_{i}}w_{ij}$ is a design-unbiased estimator of the population size $N_{i}$. The weighted direct estimates and their variances $\widehat{\text{Var}}(\hat{F}_{\alpha i}^{w})$ are computed using the direct function from the emdi R package.
Figure 1 illustrates the percent coefficient of variation (CV) of direct estimates for HCR and PG, emphasizing the increasing uncertainty in estimates for municipalities with smaller sample sizes. Statistical agencies set thresholds for the reliability of estimates based on their coefficient of variation (CV). For instance, Statistics Canada considers an estimate to be publishable if its CV is below $16.6\%$, whereas estimates with a CV exceeding $33.3\%$ are deemed unpublishable (Statistics Canada, 2015). Similarly, the UK Office for National Statistics (ONS) sets a threshold of 20% for publication suitability (ONS, 2020). If we set the reliability threshold at 20%, direct survey estimates of the 2002 LSMS dataset would be publishable for only 9 out of the 213 municipalities for HCR and just 2 for PG. This finding underscores the limitations of direct estimation methods in small domains, where high variability often renders estimates unreliable.
Figure 1: Histograms of Percent coefficient of variation of direct estimates for HCR and PG for all sampled municipalities using 2002 LSMS data.
The application of SAE methods based on the EBP significantly increases the number of municipalities for which reliable estimates can be obtained. This approach requires fitting an appropriate model to the survey data. Estimated model parameters for fixed and random effects are then combined with known population information for each municipality to predict its levels of HCR and PG. The selection of covariates is based on prior poverty assessment studies in Albania, ensuring consistency with previous methodologies. We employ the same household-level covariates as Betti, Neri and Bellini (2003) and Tzavidis et al. (2008), utilizing information available from both the survey and Census. Specifically, the following variables are considered: household size (a strong predictor of poverty), dwelling facilities (e.g., presence of TV, satellite dish, refrigerator, air conditioning, personal computer), ownership variables (e.g., homeownership, land ownership, car ownership).
To better understand the distribution of these key covariates, Table 1 presents the proportions for binary variables and the means and medians for continuous variables from both the LSMS and Census data. This provides insights into the variability of the covariates and their potential predictive power for poverty estimation. Examining these statistics helps assess the representativeness of the sample and the suitability of the selected covariates for small area estimation.
Table 1: Descriptive statistics of covariates from LSMS and Census data.
Variable LSMS Data Census Data
Median Mean/Prop Std.err. Median Mean/Prop Std.err.
Household size 4 4.28 1.76 4 4.22 1.75
homeownership 0.93 0.26 0.93 0.25
land ownership 0.50 0.50 0.50 0.50
car ownership 0.10 0.30 0.08 0.27
presence of TV 0.95 0.21 0.90 0.30
satellite dish 0.21 0.41 0.23 0.42
refrigerator 0.83 0.37 0.74 0.44
air conditioning 0.02 0.13 0.01 0.11
personal computer 0.02 0.14 0.01 0.12
The descriptive statistics highlight a strong consistency between LSMS and Census data across most covariates, reinforcing the validity of using both sources for small area estimation. Household size exhibits nearly identical distributions, with the same median and very similar mean values, suggesting that the sample adequately represents population characteristics. Similarly, ownership variables such as homeownership and land ownership display minimal differences between the two sources, indicating a high degree of reliability in their measurement. Although minor discrepancies arise in asset ownership, such as car ownership and household appliances, these differences are within an expected range and may be attributed to reporting variations between the survey and the Census. Notably, the proportion of households with a television is slightly higher in LSMS than in the Census ($0.95$ vs. $0.90$), while similar patterns are observed for refrigerator and satellite dish ownership. These variations could reflect differences in reporting behavior or coverage between the two data sources. Overall, the alignment of key covariates between the two datasets provides confidence in their joint use for poverty estimation and underscores the representativeness of the survey data in capturing relevant socioeconomic conditions.
4 Nested error linear regression model with high dimensional parameters
Unlike the traditional NER model used in Molina and Rao (2010), this paper extends the NERHDP model for small area poverty estimation, which accounts for heterogeneity across small areas in both regression coefficients and sampling variances. Specifically, let $Y_{ij}$ and $\mathbf{x}_{ij}$ denote the values of the study variable and a $p\times 1$ vector of known auxiliary variables for $j$th individual of the $i$th small area. The NERHDP model can be expressed as follows:
$Y_{ij}=\beta_{0}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{i}+\gamma_{i}+\epsilon_{ij},\quad j=1,\cdots,N_{i};\quad i=1,\cdots,m,$ (10)
where the area effects $\gamma_{i}\overset{\mathrm{iid}}{\sim}N(0,\sigma_{\gamma}^{2})$ and the errors $\epsilon_{ij}\sim N(0,\sigma_{\epsilon i}^{2})$ are independent. To represent the model at the area level, we define the following vectors and matrices for areas $i=1,\cdots,m$:
$\mathbf{y}_{i}=\underset{1\leq j\leq N_{i}}{\text{col}}(Y_{ij}),\;\bm{\epsilon}_{i}=\underset{1\leq j\leq N_{i}}{\text{col}}(\epsilon_{ij}),\;\mathbf{X}_{i}=\underset{1\leq j\leq N_{i}}{\text{col}}(\mathbf{x}^{\prime}_{ij}).$
Under this model, the response vectors $\mathbf{y}_{i}$, $i=1,\cdots,m$, are independent with $\mathbf{y}_{i}\sim N(\bm{\mu}_{i},\mathbf{V}_{i})$, where:
$\bm{\mu}_{i}=\beta_{0}\mathbf{1}_{N_{i}}+\mathbf{X}_{i}\bm{\beta}_{i}\;\;\text{and}\;\;\mathbf{V}_{i}=\sigma_{\gamma}^{2}\bm{1}_{N_{i}}\bm{1}_{N_{i}}^{\prime}+\sigma_{\epsilon i}^{2}\mathbf{I}_{N_{i}}.$
Here, $\mathbf{1}_{N_{i}}$ is an $N_{i}\times 1$ vector of ones, and $\mathbf{I}_{N_{i}}$ is the $N_{i}\times N_{i}$ identity matrix. Following with Molina and Rao (2010), for $i=1,\cdots,m$, we can decompose $\mathbf{y}_{i}$ into sample and out-of-sample elements $\mathbf{y}_{i}=(\mathbf{y}^{\prime}_{is},\mathbf{y}^{\prime}_{ir})^{\prime}$, and the corresponding decomposition on $\mathbf{X}_{i}$, $\bm{\mu}_{i}$ and $\mathbf{V}_{i}$. When the sample size $n_{i}>0$, the distribution of $\mathbf{y}_{ir}$ given the sample data is
$\mathbf{y}_{ir}|\mathbf{y}_{s}\sim N(\bm{\mu}_{ir|s},\mathbf{V}_{ir|s}),$ (11)
where
$\bm{\mu}_{ir|s}$ $=$ $\beta_{0}\mathbf{1}_{N_{i}-n_{i}}+\mathbf{X}_{ir}\bm{\beta}_{i}+\sigma_{\gamma}^{2}\bm{1}_{N_{i}-n_{i}}\bm{1}_{n_{i}}^{\prime}\mathbf{V}^{-1}_{is}(\mathbf{y}_{is}-\beta_{0}\mathbf{1}_{n_{i}}-\mathbf{X}_{is}\bm{\beta}_{i})$ (12)
$\mathbf{V}_{ir|s}$ $=$ $\sigma_{\gamma}^{2}(1-B_{i})\bm{1}_{N_{i}-n_{i}}\bm{1}_{N_{i}-n_{i}}^{\prime}+\sigma_{\epsilon i}^{2}\mathbf{I}_{N_{i}-n_{i}},$ (13)
where $\mathbf{V}_{is}=\sigma_{\gamma}^{2}1_{n_{i}}1_{n_{i}}^{\prime}+\sigma_{\epsilon i}^{2}\mathbf{I}_{n_{i}}$ and $B_{i}=\sigma_{\gamma}^{2}(\sigma_{\gamma}^{2}+\sigma_{\epsilon i}^{2}/n_{i})^{-1}$.
To avoid the computational burden of simulating multivariate normal vectors $\mathbf{y}_{ir}$ of sizes $N_{i}-n_{i},~{}i=1,\cdots,m$, the conditional distribution in (11) can be reformulated using the following model:
$\mathbf{y}_{ir}=\bm{\mu}_{ir|s}+u_{i}\mathbf{1}_{N_{i}-n_{i}}+\bm{e}_{ir},$ (14)
where $u_{i}$ represents new random effects, and $\bm{\epsilon}_{ir}$ denotes the error terms. These components are independent and satisfy the following distributions:
$u_{i}\sim N(0,\sigma_{\gamma}^{2}(1-B_{i})),\quad i=1,\cdots,m,\quad\text{and}\quad\bm{e}_{ir}\sim N(\mathbf{0}_{N_{i}-n_{i}},\sigma_{\epsilon i}^{2}\mathbf{I}_{N_{i}-n_{i}}).$
This reformulation simplifies the simulation process while maintaining the statistical properties of the original model.
As mentioned before, in practice the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, need to be replaced by suitable estimators $\hat{\bm{\phi}}_{i}=(\hat{\beta}_{0},\hat{\bm{\beta}}^{\prime}_{i},\hat{\sigma}_{\gamma}^{2},\hat{\sigma}^{2}_{\epsilon i})$. Then the variables $Y_{ij}$ are generated from the corresponding estimated normal distributions. Given the practical challenges in directly linking the survey sample to the population and under the assumption that the sampling fraction is negligible, for in-sample areas, the elements $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated from
$Y_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\hat{\sigma}_{\gamma}^{2}\mathbf{1}_{n_{i}}^{\prime}\hat{\mathbf{V}}^{-1}_{is}(\mathbf{y}_{is}-\hat{\beta}_{0}\mathbf{1}_{n_{i}}-\mathbf{X}_{is}\hat{\bm{\beta}}_{i})+u_{i}+e_{ij},$
where $u_{i}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}_{\gamma}^{2}(1-\hat{B}_{i}))$, and $e_{ij}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}^{2}_{\epsilon i})$. If an area $i$ is not sampled, then $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated by bootstrap from
$Y_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\gamma_{i}^{*}+\epsilon^{*}_{ij},$
where $\gamma_{i}^{*}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}_{\gamma}^{2})$ and $\epsilon^{*}_{ij}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}^{2}_{\epsilon i})$.
Repeat the above generation process $K$ times. Then use the formulas in (8) and (9) obtain the EBP approximation of $F_{\alpha i}$:
$\hat{F}^{EB}_{\alpha i}\approx\frac{1}{K}\sum_{k=1}^{K}\left\{\frac{1}{N_{i}}\sum_{j}^{N_{i}}h_{\alpha}(Y_{ij}^{(k)})\right\},$
for all areas of interest.
4.1 Estimation of the high dimensional parameters
In this study, we introduce a data-driven method for estimating the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, which effectively leverages data across areas to provide consistent estimators for the area-specific parameters when the tuning parameters are known. Moreover, our approach offers a significant improvement over Lahiri and Salvati (2023) by drastically reducing computational time while maintaining comparable accuracy and performance. These enhancements make our method both more efficient and practical for the analysis of large-scale datasets. The algorithm is detailed in the following steps.
  • Step 1
    Define $r_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-a_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\bm{\beta}_{i})$, where $q_{i}$ is a known scale parameter, ${\mathbf{y}}_{ls}$ is a $n_{l}\times 1$ vector of the response variable and $\mathbf{X}_{ls}$ denotes the matrix of dimension $n_{l}\times p$ individual-level covariates of the sampled units in area $l$. Obtain the area-specific regression coefficients estimates $\{(\hat{a}_{0i},\hat{\bm{\beta}}_{i}),i=1,\cdots,m\}$ via solving the following system of equations:
    $\sum_{l=1}^{m}\mathbf{X}^{\prime}_{ls(p+1)}q_{i}^{1/2}\psi_{i}(r_{l;i})=0,$ (15)
    where $\mathbf{X}_{ls(p+1)}$ is the matrix of dimension $n_{l}\times(p+1)$ including the intercept term and covariates of the sampled units in area $l$, and $\psi_{i}(r_{l;i})$ is a $n_{l}\times 1$ vector obtained from the vector of $r_{l;i}$ with its $j$th element, say $r_{lj;i}$ replaced by $\psi_{i}(r_{lj;i})$, a chosen known function of $r_{lj;i}$.
  • Remark 1
    In this section, we assume that the function $\psi_{i}$ is fully specified. For example, consider the form:
    $\psi_{i}(r)=2\psi(r)[\tau_{i}I(r>0)+(1-\tau_{i})I(r\leq 0)],-\infty<r<\infty,$ (16)
    where $\psi(r)$ is a known monotone non-decreasing function satisfying $\psi(-\infty)<\psi(0)<\psi(\infty)$, and $\tau_{i}\in T=(0,1)$ is a known tuning parameter. Notably, setting $\tau_{i}=0.5$ with the identify function for $\psi$ would lead to the standard least square estimator of the regression coefficient vector. The tuning parameter $\tau_{i}$ introduces small area-specific effects into the estimation of the regression coefficients $\bm{\beta}_{i}$ and the sampling variances $\sigma^{2}_{\epsilon i}$, enhancing the flexibility and adaptability of the model. Following Chambers and Tzavidis (2006), we assume that $\psi(\cdot)$ is the Huber influence function, with a tuning constant $c=1.345$, ensuring robust performance in the presence of outliers.
  • Step 2
    Define $\mathbf{\tilde{r}}_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-\hat{a}_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\hat{\bm{\beta}}_{i})$ and $\tilde{\mathbf{V}}_{l;i}=\sigma^{2}_{\gamma}\bm{1}_{n_{l}}\bm{1}_{n_{l}}^{\prime}+\sigma^{2}_{\epsilon i}\mathbf{I}_{n_{l}}$. Obtain $\{\hat{\sigma}^{2}_{\epsilon i},i=1,\cdots,m\}$ as a solution of the following system of estimation equations:
    $\sum_{l=1}^{m}\Bigg{[}\psi_{i}(\tilde{\mathbf{r}}_{l;i})^{\prime}\;\tilde{\mathbf{V}}_{l;i}^{-1}\;\frac{\partial\mathbf{V}_{i;i}}{\partial\sigma^{2}_{\epsilon i}}\;\tilde{\mathbf{V}}_{l;i}^{-1}\;\psi_{i}(\tilde{\bm{r}}_{l;i})\;-\text{tr}(\tilde{\mathbf{V}}_{l;i}^{-1}\;\frac{\partial\mathbf{V}_{i;i}}{\partial\sigma^{2}_{\epsilon i}})\Bigg{]}=0$ (17)
    where $i=1,\cdots,m$, and $\mathbf{V}_{i;i}=\sigma^{2}_{\gamma}\bm{1}_{n_{i}}\bm{1}_{n_{i}}^{\prime}+\sigma^{2}_{\epsilon i}\mathbf{I}_{n_{i}}$.
  • Step 3
    Define $\bar{\mathbf{r}}^{\star}=\{\bar{r}_{i}^{\star};i\in s\}^{\prime}$, where $\bar{r}_{i}^{\star}=n_{i}^{-1}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$ and $r_{ij}^{\star}=Y_{ij}-\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}$ with $\hat{\bm{\beta}}_{i}$ from Step 1. We estimate $\beta_{0}$ by taking the grand residual mean: $\hat{\beta}_{0}=1/n\sum_{i\in s}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$. Define $v_{i}^{\star}=\sigma^{2}_{\gamma}+\hat{d}_{i}$, where $\hat{d}_{i}=\hat{\sigma}^{2}_{\epsilon i}/n_{i}$ for $i\in s$. Obtain $\hat{\sigma}^{2}_{\gamma}$ as a solution of the following estimating equation:
    $\sum_{i\in s}\bigg{[}\psi^{2}\bigg{(}(v_{i}^{\star})^{-1/2}\bar{r}_{i}^{\star}\bigg{)}(v_{i}^{\star})^{-1}-(v_{i}^{\star})^{-1}w^{\star}\bigg{]}=0$ (18)
    where $w^{\star}=\text{E}[\psi^{2}(u)]$ with $u$ following a standard normal distribution.
4.2 Definition and estimation of tuning parameter $\tau_{i}$
Let the subscript $U$ denotes a finite population indicators. Define the following finite population parameters:
$Y_{ij;\tau}$ $=$ $\alpha_{0\tau;U}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{\tau;U};$
$\tau_{ij}$ $=$ $\text{argmin}_{\tau\in T}(Y_{ij;\tau}-Y_{ij})^{2},$
for $i=1,\cdots,m$; $j=1,\cdots,N_{i}$. Note that $Y_{ij;\tau}$ and $\tau_{ij}$ are unknown even for sampled units since $\alpha_{0\tau;U}$ and $\bm{\beta}_{\tau;U}$ are unknown finite population parameters. We are interested in estimating the following finite population parameter: $\tau_{i}=1/N_{i}\sum_{j=1}^{N_{i}}\tau_{ij}$.
For the sampled areas, we use the same data-driven method as Lahiri and Salvati (2023) to estimate the tuning parameters, $\tau_{i}$ with $i\in s$. Specifically, for a fine grid $\tau\in T$, we obtain a collection of fitted values for the entire sample:
$\hat{Y}_{ij;\tau}$ $=$ $\hat{\alpha}_{0\tau}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{\tau},\quad i\in s;\quad j=1,\cdots,n_{i},$
using the standard quantile or M-quantile methods, where $\hat{\alpha}_{0\tau}$ and $\hat{\bm{\beta}}_{\tau}$ are estimated intercept and regression coefficients, respectively. For each observation $Y_{ij}$ in the sample, we identify the fitted line that minimizes the prediction error, defined as the difference between $Y_{ij}$ and the predicted value from the fitted regression at $\mathbf{x}_{ij}$. Let $\hat{\tau}_{ij}$ denote the value of $\tau$ in the grid corresponding to this best-fitting line. The variability of $\hat{\tau}_{ij}$ reflects unit-level variability. Provided there are sample observations in area $i$, and a non-informative sampling method has been used to obtain them, one can consider an estimate of the area-$i$-specific tuning parameter is the sample average of $\hat{\tau}_{ij}$ for that area: $\bar{\hat{\tau}}_{i}=n_{i}^{-1}\sum_{j\in s_{i}}\hat{\tau}_{ij}$. Lahiri and Salvati (2023) also introduced an empirical linear best (ELB) predictor for $\tau_{i}$ denoted as $\hat{\tau}_{i}^{\text{ELB}}$, which enhances the performance of $\bar{\hat{\tau}}_{i}$ by leveraging data from all areas. For a detailed explanation, we refer readers to their paper.
For out-of-sample areas with no observations, we propose the following unmatched model to leverage information from observed areas, enabling the estimation of area-specific tuning parameters for completely unobserved regions:
$\text{E}(\bar{\hat{\tau}}_{i}|\tau_{i})=\tau_{i},\quad\text{Var}(\bar{\hat{\tau}}_{i}|\tau_{i})=\Delta_{i};$
$\text{E}[\text{logit}(\tau_{i})]=\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta},$
where $i=1,\ldots,m$. The variance components $\Delta_{i}$ are typically assumed to be known; however, in practice, they need to be estimated. In this study, we employ a smoothing technique to estimate $\Delta_{i}$
$\hat{\Delta}_{i}=\frac{1}{n_{i}}\sum_{i\in s}\sum_{j=1}^{n_{i}}\frac{(\hat{\tau}_{ij}-\bar{\hat{\tau}}_{i})^{2}}{n-m_{s}},\quad i\in s,$
where $m_{s}$ is the number of sampled areas. Here, $\mathbf{\bar{Z}}_{i}$ represents the population mean of relevant auxiliary variables, and $\bm{\eta}$ is the unknown vector of coefficients, estimated by $\hat{\eta}$ that minimizes the following objective function, $\mathbf{Q}(\bm{\eta})$:
$\mathbf{Q}(\bm{\eta})=\sum_{i\in s}\frac{1}{\hat{\Delta}_{i}}\left[\bar{\hat{\tau}}_{i}-\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}\right]^{2}.$
The estimated tuning parameters for the out-of-sample areas are then obtained as follows:
$\hat{\tau}_{i}=\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}},\quad i\notin s.$
Then, the area-specific model parameters, $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, can be estimated using the algorithm outlined in Section 4.1, with the tuning parameters $\tau_{i}$ replaced by their estimates, $\hat{\tau}_{i}$, for all small areas.
5 Uncertainty measures
The assessment of uncertainty of poverty indicators is crucial to analyze the quality of estimates. The MSPE of $\hat{F}^{EB}_{\alpha i}$ is a conventional measure fulfilling this goal, which is given by
$\text{MSPE}(\hat{F}^{EB}_{\alpha i})=E_{\bm{\phi}}(\hat{F}^{EB}_{\alpha i}-F_{\alpha i})^{2},$ (19)
where $E_{\bm{\phi}}$ denotes expectation with respect to the super-population model. However, the analytical approximations to the MSPE are difficult to derive when involving complex parameters such as the FGT poverty measures.
In this paper, we propose a parametric bootstrap method that can be applied to produce reasonable estimators of various uncertainty measures, not necessarily MSPE. Our bootstrap scheme builds on the method introduced by Lahiri and Salvati (2023), which demonstrated that for known $\tau_{i}$, the parametric bootstrap estimators of uncertainty measures converge in probability to the true corresponding uncertainty measures as $m$ tends to infinity, under certain regularity conditions. Our approach adapts this framework to the context of poverty mapping. The steps of the proposed bootstrap are as follows:
Step 1.
Fit model (10) to sample data and obtain estimators $\hat{\bm{\phi}}_{i}=(\hat{\beta}_{0},\hat{\bm{\beta}}^{\prime}_{i},\hat{\sigma}_{\gamma}^{2},\hat{\sigma}^{2}_{\epsilon i})$ for $i=1,\cdots,m$, using the proposed method in section 4;
Step 2.
Given $\hat{\bm{\phi}}_{i}$ from Step 1, generate $B$ parametric bootstrap populations using the following model:
$Y^{*(b)}_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+u_{i}^{*(b)}+e^{*(b)}_{ij},$ (20)
where $u_{i}^{*(b)}\sim N(0,\hat{\sigma}_{\gamma}^{2})$ and $e^{*(b)}_{ij}\sim N(0,\hat{\sigma}^{2}_{\epsilon i})$ are all independently distributed, $i=1,\cdots,m$, $j=1,\cdots,N_{i}$.
Step 3.
For each population, calculate bootstrap population parameters $F^{*(b)}_{\alpha i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}F^{*(b)}_{\alpha ij}$, where $F^{*(b)}_{\alpha ij}=h_{\alpha}(Y^{*(b)}_{ij})$, $b=1,\cdots,B$.
Step 4.
Using the same $\hat{\bm{\phi}}_{i}$ from Step 1, generate bootstrap samples with the same size as the original sample. Specifically, for each bootstrap population $b$, we have the following bootstrap sample:
$Y^{*(b)}_{ij;s}=\hat{\beta}_{0}+\mathbf{x}_{ij;s}^{\prime}\hat{\bm{\beta}}_{i}+u_{i;s}^{*(b)}+e^{*(b)}_{ij;s},i\in s,j=1,\cdots,n_{i},$ (21)
where $\{u_{i;s}^{*(b)};i\in s\}$ are selected using the same area indices as in the original sample from $\{u_{i}^{*(b)};i=1,\cdots,m\}$, which are obtained in Step 2. Here $\{e^{*(b)}_{ij;s};i\in s,j=1,\cdots,n_{i}\}$ are independently generated from $N(0,\hat{\sigma}^{2}_{\epsilon i})$. These errors are not subsets of the population errors generated in Step 2 but are instead independently drawn from the same distributions.
Step 5.
Fit model (10) to bootstrap samples and calculate the bootstrap EBPs, $\hat{F}^{EB*(b)}_{\alpha i}$, $b=1,\cdots,B$, as described in Sections 2 and 4.
Step 6.
A Monto Carlo approximation to the theoretical bootstrap estimator $\text{MSPE}_{*}(\hat{F}^{EB*(b)}_{\alpha i})=E_{\bm{\phi}^{*}}(\hat{F}^{EB*(b)}_{\alpha i}-F^{*(b)}_{\alpha i})^{2}$ of $\hat{F}^{EB}_{\alpha i}$ is calculated as
$\text{mspe}_{*}(\hat{F}^{EB}_{\alpha i})=\frac{1}{B}\sum_{b=1}^{B}(\hat{F}^{EB*(b)}_{\alpha i}-F^{*(b)}_{\alpha i})^{2}.$ (22)
The estimator $\text{mspe}_{*}(\hat{F}^{EB}_{\alpha i})$ is used to estimate $\text{MSPE}(\hat{F}^{EB}_{\alpha i})$ given in (19).
6 Design-based Monte Carlo simulations
In this section, we perform design-based simulation experiments to assess the performance of various estimators for poverty measures across repeated samples drawn from a fixed population. Specifically, we utilize the 2021 Italian dataset from the European Survey on Income and Living Conditions (EUSILC), comprising $29,798$ observations. The dataset is treated as the fixed population, with Italy’s $107$ provinces serving as the areas of interest. This setup allows us to compute the true area-level headcount ratios and poverty gaps, serving as a benchmark for evaluation:
$F_{\alpha i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}F_{\alpha ij},\quad i=1,\cdots,107,$
where
$F_{\alpha ij}=\bigg{(}\frac{z-E_{ij}}{z}\bigg{)}^{\alpha}I(E_{ij}<z),\;j=1,\cdots,N_{i}.$
Here, $E_{ij}$ denotes the equalized annual net income for each individual, and $z$ represents the fixed poverty line, which is roughly equal to 0.6 times the median of the welfare variables $E_{dj}$ for the above fixed population.
Following Molina and Rao (2010), the welfare variable $E_{ij}$ is transformed to ensure positivity by adding a fixed constant $c$, and then taking the logarithm. Specifically, the response variable is defined as $Y_{ij}=\log(E_{ij}+c)$. The auxiliary variables used in the model comprise both continuous and categorical types. The continuous variables include age, house size in square meters (MQ), household size (hsize), and total years of work experience (Years_work). Binary indicators consist of working status (Work_Status) and gender, while categorical variables include six levels of income and living conditions (cond_ilo) and marital status (Marital_status). Additionally, life satisfaction (life_satisf) is treated as a scale variable.
To explore the data, we calculate the sample variance of the response variable to understand its dispersion within each province. Additionally, we fit a regression model for the response variable using the auxiliary variables as predictors, separately for each province. This approach allows us to examine the relationships between the response and auxiliary variables while accounting for provincial differences in patterns and variability. Figure 2 shows boxplots of the distribution of residuals by province and the distribution of regression coefficients. These results indicate that the regression coefficients vary significantly across provinces, and within-area variation may also differ between provinces. Consequently, the assumptions of the traditional NER model may not hold for this dataset.
(a) Distributions of estimated regression coefficients.
(b) Distributions of the residuals by province.
Figure 2: Distributions of estimated regression coefficients (a) and residuals (b) obtained by fitting a regression model separately for each province.
In the simulation experiments, we consider the following estimators of the FGT measures for each province:
Direct estimator (DIRECT),
The EBP based on the traditional NER model from Molina and Rao (2010) (MR),
The proposed EBP based on the NERHDP model (CLS),
The ELL estimator using the algorithm from Elbers, Lanjouw and Lanjouw (2003) (ELL).
For the EBP methods, we use $K=100$ for the Monte Carlo approximation, and for the ELL method, 100 simulated censuses are generated.
Two experimental settings are considered: (1) all 107 provinces in the fixed population are sampled; (2) only 90 provinces are sampled, leaving 17 provinces with no observations. For each setting, 1000 samples are drawn using simple random sampling (SRS) without replacement within each sampled area, with the sample size set at $0.1\times N_{i}$. The performance of the above estimators is assessed using the following three criteria:
Relative Bias (RB): The average RB over all small areas. For a given area, RB is defined as the ratio of the average difference between the estimated and true small-area FGT measures to the true small-area FGT measure, averaged over simulations.
Relative Root Mean Squared Prediction Error (RRMSPE): The average RRMSPE over all small areas. For a given area, RRMSPE is the ratio of the square root of the average squared difference between the estimated and true small-area FGT measures to the true measure, averaged over simulations.
Efficiency (EFF) measured as the ratio of the RRMSPE of each estimator/predictor to the RRMSPE of the corresponding the baseline estimator. DIRECT and MR serve as the baseline estimators in experiments 1 and 2, respectively.
First, the computational efficiency of the new parameter estimation algorithm is evaluated, comparing it to the existing algorithm proposed by Lahiri and Salvati (2023). All computations are carried out on a 13th Gen Intel(R) Core(TM) i7-13700F 2.10 GHz processor. For this comparison, five samples are generated using SRS without replacement within each province, with the sample size set at $0.02\times N_{i}$, resulting in a fixed total sample size of 616. Both algorithms are applied to these five samples, and their computation times are recorded. The algorithm proposed by Lahiri and Salvati (2023) requires an average computation time of $1197.59$ seconds, while the new algorithm completes the same task in just $2.17$ seconds on average. This demonstrates a substantial improvement, as our algorithm requires only about $0.2\%$ of the computation time compared to the existing algorithm.
6.1 Experiment 1: All provinces are sampled
In this simulation experiment, we assume that all 107 provinces are included in the samples ($m_{s}=m=107$). For each province, samples are repeatedly drawn with a size equal to $0.1\times N_{i}$. The total sample size is 2979 and the province-specific sample sizes range from 2 to 180. In this scenario, all four estimators are available for all provinces.
Table 2 provides a summary of the overall performance of the four estimators, evaluated based on average RB, average RRMSPE, and EFF, with the DIRECT estimator serving as the baseline for comparison. CLS consistently outperforms MR in terms of both bias and RRMSPE, likely due to the limitations of the MR method, which assumes identical regression coefficients and variance components across the population—assumptions that may not be valid in this context. Additionally, ELL shows better performance than MR, in contrast to the findings in Molina and Rao (2010). This difference may stem from our use of the algorithm from Elbers, Lanjouw and Lanjouw (2003), which incorporates heteroscedasticity at the individual level, rather than the simplified version outlined in Molina and Rao (2010). For the poverty gap indicator, CLS outperforms in both relative bias and RRMSPE, while its performance is comparable to ELL for the headcount ratio.
Table 2: Performance of estimators/predictors of small area FGT measures when all areas are sampled; Average is over 107 areas; numbers in parenthesis are the values of the efficiency over DIRECT in terms of RRMSPE.
Predictor RRMSPE % (EFF) Relative Bias %
Headcount Ratio Poverty Gap Headcount Ratio Poverty Gap
DIRECT 48.71 (1.00) 60.43 (1.00) 4.16 5.14
MR 28.73 (0.59) 33.39 (0.55) 25.32 21.71
CLS 22.94 (0.47) 17.81 (0.29) 15.34 -2.03
ELL 22.48 (0.46) 20.25 (0.34) 15.61 8.18
Figure 3 illustrates the province-specific performance of the estimators. For the poverty gap indicator, Figure 3(a) shows that CLS outperforms MR in terms of both RRMSPE and relative bias for most provinces, with MR exhibiting a tendency toward positive bias. Similarly, Figure 3(b) highlights the performance of the four methods for estimating the headcount ratio, demonstrating that CLS achieves smaller RRMSPE and absolute bias than MR in the majority of provinces.
(a) The RRMSPE and Relative Bias of the four estimators of poverty gap for each area.
(b) The RRMSPE and Relative Bias of the four estimators of headcount ratio for each area.
Figure 3: The RRMSPE and Relative Bias of the four estimators of FGT measures, when all provinces are sampled.
6.2 Experiment 2: 17 out-of-sample provinces
In the second experiment, we assess the performance of EBP methods and the ELL method in estimating FGT measures for out-of-sample provinces, where the direct estimator is unavailable due to the absence of observations. Specifically, 90 provinces are randomly selected, and this selection is fixed to draw simple random samples (SRS) repeatedly from these provinces. Consequently, 17 provinces are consistently excluded from the sampling process, ensuring they have no observations across all simulations.
Table 3 summarizes the overall performance of the three estimators for the out-of-sample provinces. The results align with those observed in Experiment 1. Overall, CLS outperforms MR in terms of both RRMSPE and relative bias for both FGT measures. CLS also shows better performance than ELL for the poverty gap measure, while their performance is comparable for the headcount ratio.
Table 3: Performance of estimators/predictors of small area FGT measures when some areas are not sampled; Average is over 17 out-of-sample areas; numbers in parenthesis are the values of the efficiency over MR in terms of RRMSPE.
Predictor RRMSPE % (EFF) Relative Bias %
Headcount Ratio Poverty Gap Headcount Ratio Poverty Gap
MR 35.36 (1.00) 41.87 (1.00) 31.50 29.78
CLS 27.49 (0.78) 16.69 (0.40) 18.83 -0.93
ELL 26.04 (0.74) 20.55 (0.49) 19.17 10.46
Figure 4(a) illustrates the province-specific performance for the 17 out-of-sample provinces, demonstrating that the CLS model consistently outperforms both MR and ELL for the poverty gap measure in terms of RRMSPE and relative bias. Notably, CLS achieves significantly smaller RRMSPE than MR for the majority of out-of-sample areas. This superior performance can be explained by the flexibility of the CLS model, which allows regression coefficients and sampling variances to vary across areas, enabling it to better capture region-specific heterogeneity.
For the headcount ratio, Figure 4(b) shows that CLS also performs better than MR, although the margin of improvement is smaller compared to the poverty gap measure. These results highlight the capability of the CLS model to adapt to area-level differences, offering enhanced precision and reduced bias in poverty estimates.
(a) The RRMSPE and Relative Bias of the EBP estimators of poverty gap for out of sample areas.
(b) The RRMSPE and Relative Bias of the EBP estimators of headcount ratio for out of sample areas.
Figure 4: The RRMSPE and Relative Bias of the EBP estimators of FGT measures for out of sample areas.
7 Application: poverty mapping for Albania
Poverty mapping is a crucial tool for understanding the spatial distribution of poverty and guiding the implementation of poverty alleviation programs. In Albania, where the economy is predominantly rural and income is difficult to measure accurately, poverty is assessed using a consumption-based measure (Betti, Neri and Bellini, 2003). In this section, we use 2002 Living Standards Measurement Survey (LSMS) data to estimate poverty indicators for 374 municipalities in Albania. To supplement this information, we incorporate auxiliary data from the 2001 Census, which covers 726,895 households across Albania. The descriptive statistics of these datasets have been presented in Section 3, providing an overview of their key characteristics and ensuring their suitability for small area estimation.”
Design-based simulation results in Section 6 indicate that the CLS method is a promising approach due to: (i) significantly lower bias and RRMSPE compared to the MR method and (ii) greater computational efficiency relative to the existing algorithm proposed in Lahiri and Salvati (2023), improving feasibility for large-scale datasets.
To estimate poverty indicators for Albanian municipalities, we consider household consumption expenditure as the primary welfare variable, denoted as $E_{ij}$. Following Section 6, we apply a shifted logarithm transformation to obtain $Y_{ij}$ as the final response variable in the NERHDP model. The covariates, including household size, dwelling facilities, and ownership variables, are based on prior poverty assessment studies in Albania (Betti, Neri and Bellini, 2003; Tzavidis et al., 2008) and have been presented in Section 3. The poverty line is set at 4,891 Leks per month for all districts, based on Betti, Neri and Bellini (2003).
For each small area, we estimate the sample variance and fit a regression model where the response variable is the shifted logarithm transformed household consumption expenditure, with covariates previously specified from the LSMS data. Initial data exploration suggests substantial heterogeneity in both within-area variation and regression coefficients across municipalities. Figure 5 presents box plots of the distributions of estimated regression coefficients (top panel) and residual distributions by municipality (bottom panel). These results indicate that assuming identical regression coefficients and/or variance components in the MR method is unrealistic for this dataset, emphasizing the need for a more flexible modeling approach that accounts for area-specific variations.
(a) Distributions of estimated regression coefficients.
(b) Distributions of the residuals by municipality.
Figure 5: Distributions of estimated regression coefficients (a) and residuals (b) obtained by fitting a regression model separately for each municipality.
We apply the CLS method to estimate empirical best predictors (EBPs) of headcount ratios and poverty gaps by municipality, denoted as $\hat{F}_{0i}^{CLS}$ and $\hat{F}_{1i}^{CLS}$, respectively. The MSPEs of these FGT measures are estimated using the parametric bootstrap estimator $\text{mspe}_{*}(\hat{F}_{\alpha i}^{CLS})$ with $B=100$ replicates. We then compute the associated coefficients of variation (CVs) using:
$\text{CV}(\hat{F}_{\alpha i}^{CLS})=\{\text{mspe}_{*}(\hat{F}_{\alpha i}^{CLS})\}^{1/2}/\hat{F}_{\alpha i}^{CLS}.$
Following Molina and Rao (2010), we also compute direct estimators of the FGT measures $\hat{F}_{\alpha i}^{w}$ and their variance estimates $\widehat{\text{Var}}(\hat{F}_{\alpha i}^{w})$ for sampled municipalities using weighted sample means (see Section 3).
To assess the quality of our CLS estimator, we use a set of diagnostic tools based on the requirement that model-based small area estimates should align with the corresponding unbiased direct estimates, albeit more precise (Marino et al., 2019). Correlation coefficients between CLS and direct estimates for the headcount ratio and poverty gap are 0.796 and 0.802, respectively, suggesting strong agreement between the two approaches. We further evaluate coherence using the goodness-of-fit diagnostic (Brown et al., 2001):
$W=\sum_{i}^{m_{s}}\frac{(\hat{F}_{\alpha i}^{w}-\hat{F}_{\alpha i}^{CLS})^{2}}{\widehat{\text{Var}}(\hat{F}_{\alpha i}^{w})+\text{mspe}_{*}(\hat{F}_{\alpha i}^{CLS})}.$
With $W=119.59$ for headcount ratio and $W=175.75$ for poverty gap, both are below the 95th percentile of a $\chi^{2}$ distribution with 212 degrees of freedom ($\chi^{2}_{212,0.95}=177.39$), indicating that CLS estimates are coherent with direct estimates.
To assess precision gains, Table 4 presents summary statistics for CLS and direct estimates across municipalities. The results indicate that CLS estimators consistently exhibit lower CVs than direct estimators, even in municipalities with varying sample sizes. Moreover, for some municipalities where the estimated headcount ratio and poverty gap are zero, the CVs of direct estimates are unavailable, highlighting a key limitation of the direct estimation method.
Table 4: Population size, sample size, direct and CLS estimates of headcount ratios (%) and poverty gaps (%), and the associated CVs (%) of direct and CLS estimators for the Albania municipalities with sample size closest to minimum, first quartile, median, third quartile, and maximum.
Headcount Ratio Poverty Gap Municipality $N_{i}$ $n_{i}$ $\hat{F}_{0i}^{w}$ $\hat{F}_{0i}^{CLS}$ CV($\hat{F}_{0i}^{w}$) CV($\hat{F}_{0i}^{CLS}$) $\hat{F}_{1i}^{w}$ $\hat{F}_{1i}^{CLS}$ CV($\hat{F}_{1i}^{w}$) CV($\hat{F}_{1i}^{CLS}$) Hajmel $1$ 6 33.33 23.82 56.96 27.01 1.52 5.92 70.56 31.31 Orenje 1419 16 12.50 20.69 66.19 15.06 2.00 4.97 66.03 17.44 B.Curri 1488 40 20.00 24.54 32.82 8.57 5.65 5.51 43.01 9.56 Gramsh 2503 64 23.55 17.97 24.48 0.69 4.75 3.64 38.93 21.90 Tirane 89764 600 12.43 11.72 10.91 4.46 2.49 2.31 14.71 4.72
We further compare the empirical cumulative density functions (ECDFs) of the CV of the CLS estimator and the direct estimator. Figure 6 presents CV (%) values from all sampled municipalities and shows that the ECDFs for the CLS estimators, for both the headcount ratio and the poverty gap, consistently dominate those of the direct estimators, further highlighting the improved precision of CLS estimators. Additionally, we find that about $78\%$ and $89\%$ of Albanian municipalities have CV values associated with the direct estimates of the headcount ratio and poverty gap, respectively, exceeding the conventional 33% reliability threshold in small area estimation (Marino et al., 2019). In contrast, these percentages reduce to 28% and 34% when applying the proposed CLS approach.
Figure 6: CVs empirical cumulative density functions for the CLS and direct estimator.
We generate poverty maps for Albania using the direct estimation method, as shown in Figure 7. The gray areas represent regions where poverty estimates are unavailable due to insufficient sample sizes, underscoring a key limitation of the direct estimation approach in small-area analysis. In contrast, Figure 8 presents poverty maps generated using our proposed CLS method, which provides more comprehensive and reliable estimates across all municipalities. These maps reveal that several municipalities in the Bulqize district—such as Ostren, Gjorice, Shupenze, and Bulqize — exhibit the highest headcount ratio and poverty gap values, significantly exceeding those observed in other municipalities. Similarly, municipalities like Gjoaaj in the Peqin district, Katund I in the Durres district, and Arren in the Kukes district also show relatively high poverty levels, highlighting the need for targeted policy interventions to improve living conditions in these areas. In contrast, municipalities in the southern regions display substantially lower poverty levels, reflecting better economic conditions and higher living standards.
Figure 7: Municipality-level direct estimates of headcount ratios and poverty gaps in Albania.
Figure 8: Municipality-level CLS estimates of headcount ratios and poverty gaps in Albania.
These results are consistent with the findings of Betti, Neri and Bellini (2003), who used the ELL method, and those of Tzavidis et al. (2008), further validating the robustness of our approach in capturing the spatial distribution of poverty in Albania.
The CLS method offers a significant advantage over direct estimation by producing estimates for all municipalities, thereby addressing data gaps observed in the direct method. This improvement ensures a more complete understanding of poverty distribution, which is crucial for effective policy planning.
The identification of high-poverty areas, particularly in the northern and central regions, highlights the need for geographically targeted poverty alleviation programs. In contrast, the relatively lower poverty levels in the south suggest regional disparities that should be accounted for in national development strategies.
The consistency of CLS estimates across municipalities, even in those with limited survey data, reinforces the reliability of this approach in small area poverty estimation. This robustness makes it a valuable tool for informing social protection policies and resource allocation.
8 Discussion
The proposed model bridges the gap between existing approaches that either assume identity in model parameters across areas, potentially leading to misspecification, or rely on random effect specifications that demand strong distributional assumptions. By leveraging area-specific estimating equations and appropriately constructed residuals, we offer a framework for deriving empirical best prediction of complex poverty measures that captures variability across areas without the instability associated with fixed effects in small samples. On the computational side, we introduce a novel, computational efficient, and data-driven procedure for estimating model parameters, which is critical for practical applications. Additionally, we develop a new method to estimate tuning parameters for out-of-sample areas, enabling the derivation of both area-specific regression coefficients and sampling variances, even for out of sample areas.
Our simulation studies highlight the advantages of the proposed CLS method, which consistently outperforms the MR method in terms of both bias and RRMSPE, particularly when the assumptions of constant regression coefficients and sampling variances across areas are violated. These results underscore the effectiveness of our approach in handling heterogeneous and complex data structures. It is important to note that estimates for out-of-sample areas, whether derived from CLS or MR, remain synthetic due to the absence of direct observations. Future research could focus on reducing the synthetic nature of these estimates by generating pseudo-random effect estimates or incorporating other information, such as spatial correlations among areas.
The field of SAE is also evolving with the integration of non-traditional data sources, such as satellite imagery, mobile phone data, and geospatial information, alongside the use of machine learning methods to address model misspecification. For instance, Jean et al. (2016) and others illustrate how combining machine learning algorithms with SAE techniques can enhance poverty estimates, particularly in data-scarce regions. More recently, Merfeld et al. (2024) showed that integrating survey data with publicly available geospatial data using SAE can help overcome some limitations of survey data and provide more geographically disaggregated estimates of various SDG indicators at the subnational level. These advancements offer promising avenues for refining the estimation of poverty measures in contexts where traditional survey data alone are insufficient.